define the term van't hoff factor.
1
Expert's answer
2020-04-24T13:36:05-0400

Answer: The van't Hoff factor "i" is a dimensionless parameter that describes the behavior of the substance in the solution. This parameter is numerically equal to the ratio of the value of some colligative property (e.g. boiling point change "\u2206T_{bp}", melting point change "\u2206T_{mp}", osmotic pressure "\pi") of the solution of this substance to the value of the same colligative property of a non-electrolyte solution with the same concentration, when all the other conditions are the same:


"i = \frac{∆T_{bp,sln}}{∆T_{bp, nel.sln}}= \frac{∆T_{mp,sln}}{∆T_{mp, nel.sln}} = \frac{\pi_{sln}}{\pi_{ nel.sln}}"

The van't Hoff factor can be calculated using the degree of dissociation "\alpha" of the solute which formula unit contains "n" ions:


"i = 1+\alpha(n-1)."

For the strong electrolytes, the degree of dissociation "\alpha" is unity and the van't Hoff factor is equal to the number of ions in the formula unit "n" : "i=n" .